OpenStudy (anonymous):

Verify. ( I know my identities) cscx-cotx=1/cscx+cotx

OpenStudy (anonymous):

subtract

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you can maybe work with the right and left the right is \[\frac{1}{\frac{1}{b}+\frac{a}{b}}\] the left is \[\frac{1}{b}-\frac{a}{b}\]

OpenStudy (misty1212):

most of it is probably algebra

OpenStudy (misty1212):

\[\frac{1}{b}-\frac{a}{b}=\frac{1-a}{b}\]

OpenStudy (misty1212):

\[\frac{1}{\frac{1}{b}+\frac{a}{b}}\]\[=\frac{1}{\frac{1+a}{b}}\]\[=\frac{b}{1+a}\]

OpenStudy (misty1212):

so now we are at \[\frac{1-\cos(x)}{\sin(x)}\] on the the right, and \[\frac{\sin(x)}{1+\cos(x)}\] on the left

OpenStudy (misty1212):

that is a well known identity, but you can prove that without much problem if it is not

Directrix (directrix):

cscx-cotx = 1/(cscx+cotx) -------------- 1 cscx - cot x ----- * ------------ csc x + cot x csc x - cot x csc x - cot x ------------ csc^ 2x - cot^2x Recall the Pythagorean Identity: cscx - cot x ------------ (1 + cot^2x) - cot^2x 1 + cot^2x = csc^2x cscx - cot x -------------- 1 csc x - cot x

Directrix (directrix):

@JMAC1204